OpenStudy (anonymous):

Which of the following is the correct solution of -2x - 3 ≥ 21? A. x ≥ -12 B. x ≤ 12 C. x ≤ -12 D. x ≥ 12

OpenStudy (anonymous):

You need to isolate x meaning get rid of the -2 and -3 \[-2x-3+3\ge 21+3=-2x \ge24\] then divide both sides by -2 which means the inequality sign will flip \[\frac{ -2x }{ -2 }\le \frac{ 24 }{ -2 }\]

OpenStudy (anonymous):

x<12?

OpenStudy (anonymous):

Then the answer will be \[x \le \frac{ 24 }{ -2 }\] you should be able to figure out from here

OpenStudy (anonymous):

you dropped the negative before the 12 your answer should be negative

OpenStudy (anonymous):

oh alright got it, thanks

OpenStudy (anonymous):

your welcome